# Industry Expectations and Skill Demands in Quantum Software Testing

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: 7th International Workshop on Quantum Software Engineering; April 12-18, 2026; Rio de Janeiro, Braziljournalyear: 2026isbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software creation and management Software development process management
@@ 1. Introduction
Software testing is a fundamental activity in software engineering that ensures that systems behave as expected under diverse conditions. It involves systematic planning, design, execution and evaluation of tests to identify defects and improve the reliability, usability, and performance of systems (whittaker2002software). Testing is both a technical and a human-centered activity that requires a combination of analytical reasoning, tool proficiency, and collaboration among professionals with diverse roles (kassab2017software; kochhar2019practitioners). Research shows that effective testers integrate hard skills, such as programming, automation, debugging, and domain knowledge, with soft skills, such as communication, teamwork, adaptability, and critical thinking (florea2018software; florea2019skills; sanchez2020beyond; valle2023soft). As the field evolves, automation, artificial intelligence, and global collaboration are changing the nature of testing work, while the balance between technical expertise and interpersonal competence remains essential (florea2019global; gurcan2022evolution; lima2023software). This combination of technical competence and human awareness defines the foundation of professional software testing practice.
Currently, Quantum Computing introduces a new paradigm for software testing by transforming how programs are specified, executed, and verified (leite2025testing; garcia2023quantum). Quantum programs are based on principles of superposition, entanglement, and measurement, which make their behavior probabilistic and non-deterministic (de2024quantum; garcia2023quantum; paltenghi2024survey; SantaBarletta2025). Unlike classical software, the output of a quantum program cannot be fully predicted or observed, and the act of measurement collapses the system state (miranskyy2021testing; ali2023quantum). These characteristics challenge established testing processes and require the development of new techniques that account for quantum-specific features such as unitary transformations, noise, and decoherence (ali2024quantum; leite2025testing). Quantum software testing, therefore, seeks to extend traditional methods, including test input generation, mutation testing, metamorphic relations, and coverage criteria, to quantum circuits and hybrid quantum–classical systems (garcia2023quantum; paltenghi2024survey; zappin2025challenges). This emerging discipline demands both technical innovation and conceptual adaptation, creating a distinct intersection between software engineering and quantum physics.
@@
Although the body of research on quantum software testing is growing (paltenghi2024survey), there remains little understanding of how these advances are reflected in industrial practice. Organizations developing quantum technologies are rapidly building specialized teams, yet it is unclear what they expect from professionals joining these roles, especially regarding software testing experience, tool familiarity, and interpersonal abilities (zappin2025challenges; de2024quantum; leite2025testing). Understanding these expectations is key to align education, training, and professional development in Quantum Software Engineering. In this sense, this study aims to investigate what the software industry demands from practitioners engaged in quantum software testing. Our goal is to characterize the professional expectations related to software testing within the quantum software engineering industry.
@@
To achieve this, we collected and analyzed 110 publicly available job postings for positions that explicitly focus on testing or include testing responsibilities in the context of quantum software development. Based on this, we addressed the following research questions:RQ1. What testing-related activities and competencies are requested in job postings for quantum software engineers? RQ2. What hard and soft skills are expected from professionals responsible for testing quantum systems? Our contributions are threefold: (1) we provide an empirical overview of the testing expectations expressed in the quantum software job market, (2) we map the relationship between industry expectations and current research on quantum software testing, and (3) we discuss implications for education and professional training that can guide curriculum design and workforce preparation.
@@
The remainder of this paper is organized as follows. Section 2 presents the background and related work on software testing skills and quantum software testing practices. Section 3 describes our data collection and qualitative analysis procedures. Section 4 reports our findings on the testing-related demands and required skills identified in the job postings. Section 5 discusses the implications of these findings for research, education, and professional practice. Finally, Section 6 presents our conclusions and directions for future work.
@@ 2.1. Hard and Soft Skills for Software Testing
Software testing has evolved from a technical validation activity into a socio-technical practice that combines analytical reasoning, communication, and collaborative problem solving (florea2019skills; valle2023soft). Early research conceptualized testing as a rigorous and creative process that requires structured thinking and strong analytical capabilities (whittaker2002software). As software systems have become more complex and globally distributed, testing professionals are now expected to integrate technical proficiency with human-centered competencies to ensure that quality assurance activities remain effective in dynamic development environments (kassab2017software; kochhar2019practitioners). Studies have shown that the ability to design and automate tests, interpret results, and communicate findings to multidisciplinary teams has become essential to achieving reliability and efficiency in modern software projects (florea2018software; gurcan2022evolution).
Empirical investigations of industry practice demonstrate that testing requires both hard and soft skills that operate in complementary ways (valle2023soft). Hard skills refer to the technical knowledge necessary to execute, automate, and evaluate tests, including proficiency with programming languages, databases, testing frameworks, and continuous integration tools (florea2019global; florea2019skills). Soft skills encompass interpersonal and cognitive abilities such as teamwork, adaptability, communication, and critical thinking (sanchez2020beyond; valle2023soft). Surveys of practitioners reveal that testers who combine these domains perform more effectively, as technical expertise alone is insufficient to manage the ambiguity and coordination demands inherent to large-scale software projects (florea2018software; sanchez2020beyond). The ability to collaborate, negotiate priorities, and understand user perspectives has therefore become integral to the professional identity of software testers.
Recent research emphasizes that educational programs and professional training must address this duality between technical and human dimensions of testing (lima2023software; valle2023soft; florea2019skills; sanchez2020beyond; gurcan2022evolution). Analyses of curricula and job markets show persistent mismatches between academic preparation and industry expectations, with employers prioritizing candidates who demonstrate both automation competence and communication ability (lima2023software; valle2023soft). As testing increasingly relies on automation and cross-functional collaboration, organizations seek professionals who can interpret complex results, reason about system behavior, and communicate findings effectively across disciplinary boundaries (gurcan2022evolution; florea2019global). Overall, the literature establishes that the evolution of testing practice depends not only on technological innovation but also on the cultivation of hybrid skill sets that combine computational precision with human judgment (florea2018software; florea2019global; gurcan2022evolution; valle2023soft; sanchez2020beyond; kochhar2019practitioners)..
@@ 2.2. Quantum Software Testing
Quantum Computing introduces new forms of complexity into software development, requiring fundamental rethinking of how testing is defined and performed (rieffel2000introduction). Quantum programs differ from classical software because they operate on qubits rather than bits, leveraging superposition and entanglement to represent and process information (miranskyy2021testing; de2024quantum; rieffel2000introduction). These properties make the behavior of quantum systems inherently probabilistic, meaning that the same program can produce multiple valid outputs depending on measurement outcomes (rieffel2000introduction). As a result, traditional notions of correctness, determinism, and reproducibility no longer hold, creating unique challenges for verification and validation (garcia2023quantum; ali2023quantum). The inability to directly observe quantum states without altering them makes debugging complex and creates the need for new analytical techniques and tools (paltenghi2024survey).
The literature on quantum software testing has grown rapidly, outlining methods that adapt or extend classical testing principles to quantum contexts (paltenghi2024survey; Barletta2024). Researchers have proposed testing strategies that include input generation, oracle design, assertion-based verification, metamorphic testing, and mutation analysis (ali2024quantum; garcia2023quantum; paltenghi2024survey). Surveys and systematic mappings have classified these approaches according to the testing subproblems they address, including correctness validation, performance optimization, and hardware-specific analysis (paltenghi2024survey; leite2025testing). Many of these studies emphasize that quantum testing cannot rely solely on deterministic validation, requiring probabilistic models, statistical reasoning, and hybrid quantum–classical test execution environments. Despite significant conceptual progress, there remains limited empirical evidence about how these methods are implemented in practice and how they align with industrial workflows (ali2023quantum; ali2024quantum; garcia2023quantum; paltenghi2024survey; leite2025testing; zappin2025challenges).
Recent investigations of professional practice suggest that the maturity of quantum software testing is still in its formative stage. Empirical studies of developers and organizations report that most testing is performed manually, often through simulations or small-scale circuit verification (zappin2025challenges; Serrano2024). Developers rely heavily on visualization tools, print statements, and iterative debugging, while specialized quantum testing frameworks remain underutilized due to limited awareness and integration (zappin2025challenges; garcia2023quantum). Researchers have identified a clear gap between academic tool development and industrial adoption, with practitioners emphasizing the need for more scalable, user-friendly, and well-documented solutions (de2024quantum; leite2025testing). These findings indicate that while the theoretical foundation for quantum software testing is advancing, the field still lacks the professional standardization, tool support, and training infrastructure required for widespread adoption in industrial environments.
@@ 3. Method
To investigate the testing-related expectations and required skills for professionals in quantum software engineering, we conducted a systematic analysis of job postings. Job advertisements are widely used in empirical software engineering research because they offer near real-time evidence of workforce demand and detailed descriptions of professional roles (papoutsoglou2019extracting; lecca2025curious). This approach allowed us to identify how testing activities, competencies, and qualifications are represented within the emerging quantum software industry.
@@ 3.1. Data Collection
The data collection process occurred in early 2025 and followed three complementary strategies designed to ensure coverage and reliability:
@@
    Job Portals: We performed keyword-based searches on major job portals (e.g., LinkedIn, Indeed, and Glassdoor), retrieving a total of 9,818 postings. These results included a wide range of positions across the quantum technology ecosystem.
@@ 3.3. Data Analysis and Synthesis
The analysis combined descriptive and thematic approaches to provide both quantitative and qualitative insights. Descriptive statistics summarized the frequency of testing-related mentions, required skills, and professional roles, offering a general overview of the patterns observed in the job market (george2018descriptive). Thematic synthesis was then conducted to interpret the nature of testing expectations in the context of quantum software. Following guidelines for qualitative synthesis in software engineering (cruzes2011recommended), we coded textual segments describing testing activities, desired competencies, and professional attributes. The codes were iteratively refined into higher-level categories that reflected testing tasks, challenges, and skill types identified in prior literature on quantum software testing (e.g., test input generation, circuit simulation, probabilistic verification, metamorphic testing, and hybrid quantum–classical validation). This process enabled the integration of quantitative trends with conceptual understanding, highlighting how industrial job descriptions operationalize the testing concepts, challenges, and skill demands emphasized in the research literature on quantum software testing.
@@ 4.2. Quantum Software Testing Key Activities in Industry
To understand how quantum software testing is practiced in industrial settings, we compared the activities described in the literature (ali2023quantum; ali2024quantum; paltenghi2024survey; leite2025testing; zappin2025challenges; miranskyy2021testing) with the responsibilities outlined in the job postings explicitly dedicated to testing roles. The academic sources identify a range of activities that combine traditional quality assurance methods with quantum-specific validation techniques. These include automation of calibration and measurement routines, hardware-in-the-loop and device characterization, quantum error-correction and fidelity evaluation, hybrid quantum–classical interface testing, and noise-aware or probabilistic verification. Table 1 summarize these findings.
@@ 5.2. Comparing Results with the Literature
Our findings align with prior research describing the interdisciplinary and probabilistic nature of quantum software testing. Similar to earlier observations that classical testing paradigms cannot be directly transferred to quantum systems (miranskyy2021testing; ali2023quantum; garcia2023quantum), our results show that industrial practice focuses on hybrid verification strategies and calibration-driven experimentation. The prevalence of testing activities involving hardware-in-the-loop, control pulses, and fidelity evaluation confirms that organizations are prioritizing the reliability of physical systems over purely algorithmic correctness, consistent with the system-level emphasis reported by (leite2025testing) and (zappin2025challenges). These results reinforce that quantum software testing in industry remains closely tied to the operational needs of laboratory environments and device stability.
Our results also diverge from much of the current academic discourse. Whereas research on quantum testing emphasizes formalism through mutation operators, metamorphic relations, probabilistic oracles, and fault injection, industry postings rarely mention these activities. This divergence suggests a gap between the conceptual progress of research and its practical adoption. The reasons may include limited tool maturity, integration challenges, or the absence of standardized frameworks that connect academic prototypes and industrial workflows. Similar mismatches have been noted in other emerging domains of software engineering where theory advances faster than tool deployment (garcia2023quantum; ali2024quantum). In this sense, our study empirically confirms that quantum software testing remains at a formative stage marked by early experimentation rather than consolidated best practices.
Our findings also extend the literature on software testing skills. Previous studies have consistently identified the dual importance of hard and soft skills for effective testing (florea2018software; sanchez2020beyond; valle2023soft), and our results indicate that this duality persists in quantum contexts but with new interpretations. Technical skills now encompass quantum-specific programming and control, while soft skills expand to scientific collaboration and cross-domain communication. This aligns with recent observations that the success of quantum software projects depends on professionals who can mediate between software abstractions and physical experimentation (zappin2025challenges; leite2025testing). By empirically mapping these competencies, our results contribute to the ongoing discussion about what professional preparation is required for the emerging field of quantum software engineering.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: 17h International Workshop on Quantum Software Engineering ; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 17h International Workshop on Quantum Software Engineering (Q-SE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3786150.3788614isbn: 979-8-4007-2383-4/2026/04ccs: Software and its engineering Software creation and management Software development process management
@@ 1. Introduction
Software testing is a fundamental software engineering activity that plans, designs, executes, and evaluates tests to uncover defects and improve system reliability, usability, and performance (Whittaker, 2002). It is also a socio-technical practice, requiring analytical reasoning and tool proficiency alongside collaboration across roles (Kassab et al., 2017; Kochhar et al., 2019; Pardo et al., 2010). Effective testers therefore combine hard skills (e.g., programming, automation, debugging, and domain knowledge) with soft skills (e.g., communication, teamwork, adaptability, and critical thinking), while ongoing trends such as automation, AI, and global collaboration continue to reshape testing work without eliminating the need for this balance (Florea and Stray, 2018, 2019b; Sánchez-Gordón et al., 2020; Valle et al., 2023; Florea and Stray, 2019a; Gurcan et al., 2022; Lima et al., 2023).
Quantum computing reshapes software testing by altering how programs are specified, executed, and verified (Leite Ramalho et al., 2025; García de la Barrera et al., 2023; Murillo et al., 2025). Quantum behavior is inherently probabilistic and only indirectly observable due to measurement-induced collapse, and it is further affected by noise and decoherence (De Stefano et al., 2024; Paltenghi and Pradel, 2024; Santa Barletta et al., 2025; Miranskyy et al., 2021; Ali and Yue, 2023; García de la Barrera et al., 2023). These constraints challenge classical testing assumptions and motivate quantum-specific techniques (e.g., for unitary transformations) and adaptations of test generation, mutation testing, metamorphic relations, and coverage criteria to quantum circuits and hybrid quantum–classical systems (Ali, 2024; Leite Ramalho et al., 2025; Ali, 2024; García de la Barrera et al., 2023; Paltenghi and Pradel, 2024; Zappin et al., 2025).
@@
Although research on quantum software testing is expanding, there is still limited evidence on how these advances translate into industrial practice (Paltenghi and Pradel, 2024). As organizations rapidly build quantum teams, their expectations for testing-related roles—spanning prior testing experience, tool familiarity, and interpersonal competencies—remain unclear (Zappin et al., 2025; De Stefano et al., 2024; Leite Ramalho et al., 2025). To help align education, training, and professional development in Quantum Software Engineering, this study investigates industry demand and characterizes professional expectations for quantum software testing roles.
@@
To achieve this, we collected and analyzed 110 publicly available job postings [^1] Data available at https://figshare.com/s/6ac4583b0ec96e545c4d for positions that explicitly focus on testing or include testing responsibilities in the context of quantum software development. Based on this, we addressed the following research questions: RQ1. What testing-related activities and competencies are requested in job postings for quantum software engineers? RQ2. What hard and soft skills are expected from professionals responsible for testing quantum systems? Our contributions are threefold: (1) we provide an empirical overview of the testing expectations expressed in the quantum software job market, (2) we map the relationship between industry expectations and current research on quantum software testing, and (3) we discuss implications for education and professional training that can guide curriculum design and workforce preparation.
@@
The remainder of this paper is organized as follows. Section 2 reviews software testing skills and quantum software testing. Section 3 describes the data collection and qualitative analysis. Section 4 presents the testing-related demands and skills identified in the job postings, and Section 5 discusses implications for research, education, and practice. Section 6 concludes and outlines future work.
@@ 2.1. Hard and Soft Skills for Software Testing
Early work already framed testing as a rigorous yet creative activity requiring structured thinking (Whittaker, 2002). As software systems and development organizations have grown more complex and distributed, testers are increasingly expected to combine technical proficiency with human-centered competencies to sustain effective quality assurance in dynamic environments (Florea and Stray, 2019b; Valle et al., 2023; Kassab et al., 2017; Kochhar et al., 2019).
Empirical studies show that effective testing combines complementary hard and soft skills (Valle et al., 2023). Hard skills include the technical knowledge to implement, automate, and evaluate tests (e.g., programming, test frameworks, and CI tooling) (Florea and Stray, 2019a, b). Soft skills cover interpersonal and cognitive abilities—such as communication, teamwork, and critical thinking—that help testers navigate ambiguity and coordination in large projects (Sánchez-Gordón et al., 2020; Florea and Stray, 2018; Valle et al., 2023). Practitioner surveys suggest that technical expertise alone is insufficient; collaboration and priority negotiation are therefore central to the tester role (Florea and Stray, 2018; Sánchez-Gordón et al., 2020).
Educational programs and professional training increasingly need to address the dual technical and human dimensions of testing, as curricula and job-market studies report persistent mismatches between academic preparation and industry expectations (Lima et al., 2023; Valle et al., 2023; Florea and Stray, 2019b; Sánchez-Gordón et al., 2020; Gurcan et al., 2022). Employers therefore prioritize hybrid profiles that combine automation competence with the ability to interpret results, reason about system behavior, and communicate effectively in cross-functional settings (Lima et al., 2023; Valle et al., 2023; Gurcan et al., 2022; Florea and Stray, 2019a). Overall, the literature suggests that advances in testing depend not only on technological innovation but also on cultivating such blended skill sets that integrate computational precision with human judgment (Florea and Stray, 2018, 2019a; Gurcan et al., 2022; Valle et al., 2023; Sánchez-Gordón et al., 2020; Kochhar et al., 2019).
@@ 2.2. Quantum Software Testing
Quantum computing introduces testing challenges that differ from classical settings because program behavior is probabilistic and internal states are only indirectly observable, which weakens traditional assumptions about correctness, determinism, and reproducibility (Rieffel and Polak, 2000; Miranskyy et al., 2021; De Stefano et al., 2024; García de la Barrera et al., 2023; Ali and Yue, 2023). As a result, debugging and verification often require new analytical techniques and tool support tailored to quantum execution and measurement (Paltenghi and Pradel, 2024; Rieffel and Polak, 2000).
The literature on quantum software testing has expanded rapidly, proposing methods that adapt classical testing principles to probabilistic quantum behavior, including oracle- and assertion-based verification, metamorphic relations, and mutation analysis (Paltenghi and Pradel, 2024; Barletta et al., 2024; Ali, 2024; García de la Barrera et al., 2023). Surveys and mappings organize these techniques by the subproblems they target (e.g., correctness, performance, and hardware-related analysis) and commonly highlight the need for statistical reasoning and hybrid quantum–classical execution (Paltenghi and Pradel, 2024; Leite Ramalho et al., 2025). Despite this conceptual progress, empirical evidence remains limited on how such methods are implemented in practice and aligned with industrial workflows (Ali and Yue, 2023; Ali, 2024; García de la Barrera et al., 2023; Paltenghi and Pradel, 2024; Leite Ramalho et al., 2025; Zappin et al., 2025).
Recent evidence suggests that quantum software testing remains formative: empirical studies report that testing is largely manual and often relies on simulation or small-scale circuit checks, supported by ad-hoc debugging and visualization (Zappin et al., 2025; Serrano et al., 2024). Specialized testing frameworks are still underutilized due to limited awareness and integration challenges, reinforcing a gap between academic tooling and industrial adoption (Zappin et al., 2025; García de la Barrera et al., 2023). Practitioners therefore highlight the need for scalable, usable, and well-documented solutions, alongside clearer professional standardization and training pathways to enable broader industrial uptake (De Stefano et al., 2024; Leite Ramalho et al., 2025).
@@ 3. Method
To investigate the testing-related expectations and required skills for professionals in quantum software engineering, we conducted a systematic analysis of job postings. Job advertisements are widely used in empirical software engineering research because they offer near real-time evidence of workforce demand and detailed descriptions of professional roles (Papoutsoglou et al., 2019; Leça and Santos, 2025; Pardo et al., 2011). This approach allowed us to identify how testing activities, competencies, and qualifications are represented within the emerging quantum software industry.
@@ 3.1. Data Collection
The data collection took place between August and October 2025, and followed three complementary strategies designed to ensure coverage and reliability:
@@
    Job Portals: We performed keyword-based searches on major job portals (LinkedIn, Indeed, Glassdoor, Zip Recruiter, Jooble, Flexjobs, JobLeads, Quantum Jobs US and Workday), retrieving a total of 9,818 postings. These results included a wide range of positions across the quantum technology ecosystem. We used ”Quantum” as the main keyword, and complemented the search with specific technologies keywords (Qiskit, Cirq, Q#, QDK, PennyLane, QuTiP, D-Wave, OceanSDK, Quill, PyQuill, OpenQuasm).
@@ 3.3. Data Analysis and Synthesis
The analysis combined descriptive and thematic approaches to provide both quantitative and qualitative insights. Descriptive statistics summarized the frequency of testing-related mentions, required skills, and professional roles, offering a general overview of the patterns observed in the job market (George and Mallery, 2018). Thematic synthesis was then conducted to interpret the nature of testing expectations in the context of quantum software. Following guidelines for qualitative synthesis in software engineering (Cruzes and Dyba, 2011), we coded textual segments describing testing activities, desired competencies, and professional attributes. The codes were iteratively refined into higher-level categories that reflected testing tasks, challenges, and skill types identified in prior literature on quantum software testing (e.g., test input generation, circuit simulation, probabilistic verification, metamorphic testing, and hybrid quantum–classical validation). This process enabled the integration of quantitative trends with conceptual understanding, highlighting how industrial job descriptions operationalize the testing concepts, challenges, and skill demands emphasized in the research literature on quantum software testing.
@@ 4.2. Quantum Software Testing Key Activities in Industry
To understand how quantum software testing is practiced in industrial settings, we compared the activities described in the literature (Ali and Yue, 2023; Ali, 2024; Paltenghi and Pradel, 2024; Leite Ramalho et al., 2025; Zappin et al., 2025; Miranskyy et al., 2021; Patón-Romero2019180) with the responsibilities outlined in the job postings explicitly dedicated to testing roles. The academic sources identify a range of activities that combine traditional quality assurance methods with quantum-specific validation techniques. These include automation of calibration and measurement routines, hardware-in-the-loop and device characterization, quantum error-correction and fidelity evaluation, hybrid quantum–classical interface testing, and noise-aware or probabilistic verification. Table 1 summarize these findings.
@@ 5.2. Comparing Results with the Literature
Our findings align with prior research describing the interdisciplinary and probabilistic nature of quantum software testing. Similar to earlier observations that classical testing paradigms cannot be directly transferred to quantum systems (Miranskyy et al., 2021; Ali and Yue, 2023; García de la Barrera et al., 2023), our results show that industrial practice focuses on hybrid verification strategies and calibration-driven experimentation. The prevalence of testing activities involving hardware-in-the-loop, control pulses, and fidelity evaluation confirms that organizations are prioritizing the reliability of physical systems over purely algorithmic correctness, consistent with the system-level emphasis reported by (Leite Ramalho et al., 2025) and (Zappin et al., 2025). These results reinforce that quantum software testing in industry remains closely tied to the operational needs of laboratory environments and device stability.
Our results also diverge from much of the current academic discourse. Whereas research on quantum testing emphasizes formalism through mutation operators, metamorphic relations, probabilistic oracles, and fault injection, industry postings rarely mention these activities. This divergence suggests a gap between the conceptual progress of research and its practical adoption. The reasons may include limited tool maturity, integration challenges, or the absence of standardized frameworks that connect academic prototypes and industrial workflows. Similar mismatches have been noted in other emerging domains of software engineering where theory advances faster than tool deployment (García de la Barrera et al., 2023; Ali, 2024). In this sense, our study empirically confirms that quantum software testing remains at a formative stage marked by early experimentation rather than consolidated best practices.
Our findings also extend the literature on software testing skills. Previous studies have consistently identified the dual importance of hard and soft skills for effective testing (Florea and Stray, 2018; Sánchez-Gordón et al., 2020; Valle et al., 2023), and our results indicate that this duality persists in quantum contexts but with new interpretations. Technical skills now encompass quantum-specific programming and control, while soft skills expand to scientific collaboration and cross-domain communication. This aligns with recent observations that the success of quantum software projects depends on professionals who can mediate between software abstractions and physical experimentation (Zappin et al., 2025; Leite Ramalho et al., 2025). By empirically mapping these competencies, our results contribute to the ongoing discussion about what professional preparation is required for the emerging field of quantum software engineering.